# Reflection on Data Storytelling Tools in the Generative AI Era from the Human-AI Collaboration Perspective

### v1

@@
Abstract.
Human-AI collaborative tools attract attentions from the data storytelling community to lower the barrier of expertise and streamline the workflow. The recent advance in large-scale generative AI techniques, e.g., large language models (LLMs) and text-to-image models, has the potential to enhance data storytelling with their power in visual and narration generation. After two years since these techniques were publicly available, it is important to reflect our progress of applying them and have an outlook for future opportunities. To achieve the goal, we compare the collaboration patterns of the latest tools with those of earlier ones using a dedicated framework for understanding human-AI collaboration in data storytelling. Through comparison, we identify persistent collaboration patterns, e.g., human-creator + AI-assistant, and emerging ones, e.g., AI-creator + human-reviewer. The benefits of these AI techniques and other implications to human-AI collaboration are also revealed. We further propose future directions to hopefully ignite innovations.
A browser of our collected tools is available at https://human-ai.notion.site/data-storytelling-dec-2024.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visualization
1. Introduction
@@
Data storytelling is considered as one of the major research directions in visualization research (Kosara and Mackinlay, 2013). To lower the barrier of telling appealing and effective stories, researchers have spent considerable efforts to build AI-powered tools to facilitate their creation and communication with different strategies of human-AI collaboration (Li et al., 2024b). In these tools, AI collaborators are often powered by heuristic-based methods (Wang et al., 2019), traditional machine learning models (Fu et al., 2019), or smaller-scale deep learning models (Qian et al., 2020).
@@
Compared to the previous techniques for AI collaborators, the recently emerging large-scale generative AI models, including the text-to-image models (Yang et al., 2023) and large language models (LLMs) (Zhao et al., 2023), can achieve better performance on various data storytelling-related tasks, such as data analysis (Inala et al., 2024) and text generation (Zheng et al., 2023), and enhance the communication between humans and AI with conversations. These advantages indicate their potentials to be game-changers in the research direction of human-AI collaboration for data storytelling, including improving the experience of collaborating with AI and diversifying the collaboration patterns between humans and AI (Li et al., 2024b). After two years of the public access of these models, it is a critical time point to reflect how this research discipline progresses in the new era of large-scale generative AI models and identify future opportunities. To achieve the goal, it is essential not only to focus on how these generative AI techniques are applied in existing tools, as explored in a previous survey (He et al., 2024), but more importantly, to compare the human-AI collaboration patterns in the latest tools in the generative AI era with those in earlier ones. Only through this comparison can we understand the shift in human-AI collaboration paradigms, identify the value of these powerful techniques in enhancing human-AI collaboration, and propose future research directions.
To perform the comparison, the foundation is a fine-grained and technique-agnostic framework to delineate human-AI collaboration patterns in the tools both before and in the era of generative AI. We notice that the framework by Li et al. (Li et al., 2024b) can fulfill the requirement. The framework breaks down the human-AI collaboration in data storytelling tools to two dimensions, the stages in the workflow (e.g., planning and implementation) and the roles of human and AI collaborators (e.g., creator and optimizer), which are both disentangled from the AI techniques applied in tools. The authors further leveraged the framework to conduct a comprehensive survey of related tools before mid-2023. Based on the framework and previously surveyed tools, our work reviews the latest 27 data storytelling tools and compares their human-AI collaboration patterns with those of earlier tools before the public availability of these large-scale AI models.
@@
Through the comparison, we find patterns that remain popular in the new era, such as the mode where AI assists human creators, and emerging patterns, such as human as reviewers to AI-created content. Then, we discuss their implications to human-AI collaboration, such as the advantages of these models, the necessity and suitable scenarios of both human-AI collaborative and fully automatic tools, and noteworthy research opportunities, such as the dependency of roles in different stages. By reflecting on the progress so far and exploring future directions, we hope this paper serves as a cornerstone for investigating the collaboration between humans and large-scale generative AI.
2. Related Surveys
@@
To review data storytelling tools, there were several previous surveys for different focuses. The first group of surveys (Tong et al., 2018; Zhao and Elmqvist, 2023; Ren et al., 2023) reviews tools from the aspect of how they support narrative techniques, such as data story structures (Tong et al., 2018) and components (Zhao and Elmqvist, 2023; Shen et al., 2025). The second group focuses on the application of automation techniques, such as AI, in data storytelling tools. Schröder et al. (Schröder et al., 2023) only grouped AI-powered automatic tools based on their functions and genres. Chen et al. (Chen et al., 2023a) further categorized AI-powered tools into manual, mixed-initiative, and automatic types. Li et al. (Li et al., 2024b) proposed a more fine-grained framework to explicitly understand the collaboration mechanism between human and AI through their collaboration stages and roles. Recently, motivated by the advancement of generative AI models, He et al. (He et al., 2024) reviewed their usage in tools based on tasks, such as insight identification and narration generation.
Under the second category, our paper aims to identify the insights about how human-AI collaboration in data storytelling tools is advanced in the era of generative AI models. It serves as a sequel to the work by Li et al. (Li et al., 2024b) to refresh our understanding through comparing the human-AI collaboration patterns of the latest tools with the patterns of earlier tools in their work. Moreover, unlike the survey by He et al. (He et al., 2024) and others at the intersection of large-scale generative AI models and visualization (Ye et al., 2024; Yang et al., 2024), we focus less on the technical perspective of applying these models but emphasize the advancement of higher-level human-AI collaboration patterns of data storytelling tools.
3. Method and Overview of Tools
@@
Following Li et al. (Li et al., 2024b), we review the recent data storytelling tools after June 2023 and compare their collaboration patterns with those of earlier ones. This section introduces our detailed method and an overview of tools. For brevity, we denote the tools after June 2023 as the latest tools in the era of generative AI and the tools before as earlier tools.
3.1. Tool Collection
In our paper, we focus on the tools in the academia since their designs are often described with more details. To collect the corpus, we searched tools between June 2023 and Dec. 2024 in related premier human-computer interaction venues, including CHI, UIST, and IUI, and visualization venues, including VIS, PacificVis, EuroVis, and TVCG, with keywords “narrative visualization” and “data storytelling”. Finally, we collected 27 tools, as shown in Table 1.
@@ 3.1. Tool Collection
After a careful consideration, we decided to compare human-AI collaboration patterns of the latest tools after June 2023 with those of earlier tools to study the research trend of human-AI collaborative tool in the era of generative AI. First, the dates of public access to two representative large-scale generative AI techniques, ChatGPT and Stable Diffusion, are Nov. 30, 2022 (OpenAI, 2022) and Aug. 22, 2022 (Stability.ai, 2023), respectively. Considering the delay led by tool development and paper publishing, the research tools to apply them and other similar techniques might only be available in mid-2023. Then, we examined the corpus of Li et al. (Li et al., 2024b), which includes tools published before June 2023. Their corpus do not have extensive application of these large-scale generative models, while the authors have discussed the potentials to apply them. It also supports our decision. Our surveyed tools further verify the decision: there were pioneering ones that apply text-to-image models (Xiao et al., 2023) and LLMs (Sultanum and Srinivasan, 2023; Shen et al., 2023; Lin et al., 2023) published in the VIS 2023 conference in Oct. 2023.
@@
Table 1. This table presents the human-AI collaboration patterns of our collected tools with the framework of stages and roles by Li et al. (Li et al., 2024b) and the applied generative AI techniques. The collaborators’ roles are shown in abbreviation: H-C, H-A, H-O, H-R are the creator, assistant, optimizer, and reviewer roles performed by humans; A-C, A-A, A-O, A-R are the corresponding roles by AI. The techniques are denoted with the icons: * refers to LLMs, such as GPT-4 (OpenAI, 2023) and # refers to text-to-image models, such as diffusion models (Yang et al., 2023). The tools are sorted according to their covered stages and collaboration patterns.
[…]
This table includes six columns: the publication years of tools, the tool name, and the collaboration patterns in four stages: analysis, planning, implementation, and communication. There are in total 27 rows, where each row records one tool. The tools are sorted by their covered stages and collaboration patterns. Year Tool Analysis Planning Implementation Communication 2023 SwimFlow (Yao et al., 2023) N/A N/A H-C N/A 2024 Epigraphics (Zhou et al., 2024a) N/A N/A H-C A-A ∗# N/A 2024 AnalogyMate (Chen et al., 2024) N/A N/A H-C A-A ∗# N/A 2024 Emoji Encoder (Brehmer et al., 2024) N/A N/A H-C A-A N/A 2024 CAST+ (Shen et al., 2024c) N/A N/A H-C A-A N/A 2024 Cai et al. (Cai et al., 2024) N/A N/A H-C A-A N/A 2023 DataTales (Sultanum and Srinivasan, 2023) N/A N/A A-C H-O N/A 2024 VisTellAR (Tong et al., 2024) N/A H-C H-C A-A N/A 2024 WonderFlow (Wang et al., 2024b) N/A H-C H-C A-A N/A 2024 VisConductor (Femi-Gege et al., 2024) N/A H-C H-C H-C A-A 2023 Data Player (Shen et al., 2023) N/A H-C A-C N/A 2024 Data Playwright (Shen et al., 2024b) N/A H-C A-C H-O N/A 2023 ChartSpark (Xiao et al., 2023) H-C A-A N/A H-C A-A # N/A 2023 InkSight (Lin et al., 2023) H-C N/A A-C H-O N/A 2024 OutlineSpark (Wang et al., 2024a) H-C H-C A-A A-C H-O N/A 2024 LEVA (Zhao et al., 2024) H-C A-A H-C A-C N/A 2024 CoInsight (Li et al., 2024a) H-C A-A H-C A-A A-C N/A 2024 Data Director (Shen et al., 2024a) A-C A-C A-C N/A 2024 Shi et al. (Shi et al., 2024) A-C A-C A-C N/A 2024 Live Charts (Ying et al., 2024) A-C A-C A-C A-C 2024 Sportify (Lee et al., 2024) A-C A-C A-C A-C 2024 AiCommentator (Andrews et al., 2024) A-C A-C A-C A-C 2023 Calliope-Net (Chen et al., 2023b) A-C H-O A-C H-O A-C H-O N/A 2024 DG Comics (Kim et al., 2024) A-C H-O A-C H-O A-C H-O N/A 2024 SNIL (Cheng et al., 2024) A-C H-O A-C H-O A-C H-O N/A 2023 Socrates (Li et al., 2024a) A-C H-R A-C H-R A-C N/A 2024 Han and Isaacs (Han and Isaacs, 2024) N/A N/A N/A H-C A-A
@@
Figure 1. The change of human-AI collaboration pattern frequencies between the latest and earlier tools in four stages. In a subfigure, the color of blocks and numbers in the heatmap encodes the change of pattern frequencies in the stage. A pattern is represented by a human role on the horizontal axis and an AI role on the vertical axis. Green and positive numbers mean that the pattern is more frequent in the latest tools while red and negative numbers indicate that the pattern is less frequent. A number in brown means that the pattern is newly identified in the latest tools. An empty white cell means that the pattern has not emerged in any tools. Two bar charts on the right and the top of the heatmap show the frequency change of one role between the latest and earlier tools.
[…]
This figure shows the change of collaboration patterns in four stages. The four subfigures show the pattern changes in four stages. In the analysis stage, the pattern of human-creator only decreases 23%, human-creator and AI-assistant increases 3%, AI-creator and human-assistant decreases 6%, AI-creator and human-optimizer decreases 4%, AI-creator and human-reviewer, a new pattern, increases 7%, and AI-creator only increases 23%. In the planning stage, the pattern of human-creator only decreases 16%, human-creator and AI-assistant increases 1%, human-creator and AI-optimizer decreases 7%, AI-creator and human-optimizer decreases 10%, AI-creator and human-reviewer, a new pattern, increases 6%, and AI-creator only increases 26%. In the implementation stage, the pattern of human-creator only decreases 15%, human-creator and AI-assistant increases 10%, AI-creator and human-assistant decreases 5%, AI-creator and human-optimizer decreases 8%, AI-assistant only decreases 2%, AI-reviewer only decreases 3%, and AI-creator only increases 23%. In the communication stage, the pattern of human-creator only decreases 60%, human-creator and AI-assistant decrease 7%, and AI-creator only increases 67%. The other patterns that are not mentioned above have not been implemented in existing tools.
3.2. Pattern Coding
@@
After collecting all tools, we coded their human-AI collaboration patterns with the framework proposed by Li et al. (Li et al., 2024b). The framework has two dimensions: the roles of human and AI collaborators (i.e., human-creator and AI-creator to complete the majority of tasks in each stage, human-assistant and AI-assistant to help creators on designated tasks in a stage, human-optimizer and AI-optimizer to improve the quality of the output of one task or the entire stage autonomously, and human-reviewer and AI-reviewer to provide suggestions for the output from a task or stage) and their collaboration stages (i.e., analysis stage to derive data insights and knowledge from datasets, planning stage to decide how to convey analytical findings through a data story, implementation stage to create data story content following the plan, and communication stage to share data stories directly with audiences). One author coded all tools after understanding the coding principles and the coded tools by Li et al. comprehensively. During the coding process, the results were also calibrated with previously coded tools as references. The final results are presented in Table 1 and an online browser: https://human-ai.notion.site/data-storytelling-dec-2024. To facilitate comparison, we computed and visualized the changes in the relative frequencies [^1] All frequencies refer to relative frequencies in this paper to accommodate the different numbers of tools in each stage in earlier and the latest tools. of collaboration patterns in Figure 1(d). The frequency of a pattern $p$ in a stage $s$ is computed as $f=Occurrence_{p\textrm{ in }s}/NumTool_{s}$ and the frequency change between the latest tools and earlier tools is $f_{latest}-f_{earlier}$. Therefore, a positive change in Figure 1(d) means that the pattern is gaining popular while a negative one shows the frequency is declining. Figures 3(d) and 4(d) show the frequencies of earlier and the latest tools.
We decided to apply the framework for a series of reasons. First, the framework provides a more fine-grained approach to distinguish the collaboration patterns of human and AI in the storytelling workflow than others (e.g.(Chen et al., 2023a)). Furthermore, the framework was summarized based on an empirical interview study (Li et al., 2023a) that envisions the potential usage and challenges of AI collaborators after these large-scale AI models were released. Checking the alignment between existing tools and user expectations can help us reflect and identify opportunities. Another advantage is that the framework is technique-agnostic, which allows us to compare the collaboration patterns without interference led by the differences in techniques. Lastly, earlier tools examined with the framework by Li et al. (Li et al., 2024b) can be easily re-used for comparison. The results were verified through peer review and thus provide solid foundation for our research.
3.3. Overview of Tools
@@
Combining the results of ours and Li et al. (Li et al., 2024b), we notice a considerable growth in the number of data storytelling tools in 2024 (Figure 2). Compared to the 11 tools published in 2023, the 20 tools in 2024 nearly double the previous number, reaching a historic high. Furthermore, 16 out of the 27 latest tools apply generative AI models, where 13 tools use LLMs merely, one tool uses text-to-image models merely, and two tools use both (Table 1). These models are most frequently used in the implementation stage (15/27), following up by the analysis (5/27), the planning (4/27), and the communication stages (3/27). The findings reveal that large-scale generative AI models has been widely applied across the whole storytelling workflow, reinforcing our motivation of understanding changes in the generative AI era.
4. Findings
Based on Figure 1(d), this section introduces insights about the change of collaboration patterns, including the continuously popular patterns (Section 4.1), the new patterns (Section 4.2), and the patterns that attract more attention (Section 4.3).
4.1. What pattern remains popular?

### v2

@@
Abstract
Human-AI collaborative tools attract attentions from the data storytelling community to lower the expertise barrier and streamline the workflow. The recent advance in large-scale generative AI techniques, e.g., large language models (LLMs) and text-to-image models, has the potential to enhance data storytelling with their power in visual and narration generation. After two years since these techniques were publicly available, it is important to reflect our progress of applying them and have an outlook for future opportunities. To achieve the goal, we compare the collaboration patterns of the latest tools with those of earlier ones using a dedicated framework for understanding human-AI collaboration in data storytelling. Through comparison, we identify consistently widely studied patterns, e.g., human-creator + AI-assistant, and newly explored or emerging ones, e.g., AI-creator + human-reviewer. The benefits of these AI techniques and implications to human-AI collaboration are also revealed. We further propose future directions to hopefully ignite innovations.
keywords:
@@ Abstract
1 Introduction
@@
Data storytelling is considered one of the major research directions in visualization research. To facilitate telling appealing and effective stories, researchers have spent considerable efforts to build AI-powered tools with different strategies of human-AI collaboration [21]. In these tools, AI collaborators are often powered by heuristic-based methods, traditional machine learning models, or smaller-scale deep learning models [39].
@@
Compared to previous techniques used for AI collaborators, recently emerging large-scale generative AI models, including the text-to-image models and large language models (LLMs), can achieve better performance on various data storytelling-related tasks, such as data analysis and text generation, and enhance human-AI communication through conversational interactions. These advantages highlight their potentials to be game-changers in the research area of human-AI collaboration for data storytelling, including improving the experience of working with AI and diversifying the collaboration patterns between humans and AI [21]. After two years of the public release of these large-scale generative models, this is a critical moment to reflect on how this research discipline progresses, identify the benefits of those powerful techniques comprehensively, and propose future research opportunities.
@@
To achieve the goal, we compared the human-AI collaboration patterns of 27 latest tools in the generative AI era with those of earlier tools, based on the roles of human and AI and the stages where they collaborate in the data storytelling workflow [21]. Through the comparison, we find patterns that are still widely studied in the new era, such as the mode where AI assists human creators, and newly explored or emerging patterns, such as humans acting as reviewers of AI-generated content. We then discuss the implications for human-AI collaboration, such as the advantages of these large-scale generative AI models, the necessity and suitable scenarios for human-AI collaborative or fully automatic tools, and noteworthy research opportunities, such as the dependency of roles in different stages. Through discussions on the progress and future opportunities, we hope this paper can serve as a cornerstone for investigating the collaboration between humans and large-scale generative AI.
2 Related Surveys
@@
To review the progress of data storytelling tools, the first group of surveys focuses on how tools support narrative techniques, such as data story structures [35] and components [29]. The second group focuses on the application of automation techniques, such as AI, in data storytelling tools. Chen et al. [4] categorized AI-powered tools into manual, mixed-initiative, and automatic types. Li et al. [21] proposed a fine-grained framework to explicitly characterize the collaboration between human and AI through collaboration stages and roles. Recently, He et al. [12] reviewed generative AI model usage in tools based on tasks, such as insight identification. However, its taxonomy does not consider how humans work with AI models. Under the second category, our paper aims to identify insights about how the research in human-AI collaboration for data storytelling is advanced in the era of generative AI models. As a sequel to the work by Li et al. [21], it refreshes our understanding through comparing the human-AI collaboration patterns of the latest tools with those of earlier tools and summarizes future opportunities.
3 Method
@@
This section introduces our research method of collecting tools and analyzing the collaboration patterns. For brevity, we denote the tools after Jun. 2023 as the latest tools in the generative AI era and the tools before as earlier tools.
3.1 Tool Collection
To collect the corpus, we searched for papers published between Jun. 2023 and Dec. 2024 in related premier HCI venues, including CHI, UIST, and IUI, and visualization venues, including VIS, PacificVis, EuroVis, and TVCG, with keywords “narrative visualization” and “data storytelling”. We also supplemented the corpus with papers that were reviewed in a latest related survey [29]. Following Li et al. [21], we then filtered out the papers that are not for storytelling purpose with infomation visualization or do not describe tools. Finally, we collected 27 tools, as shown in Table 1. As a preliminary investigation, this paper focuses on the tools in the academia since these tools are often described with more details, such as design rationales and usage scenarios, making it easier to identify the high-level collaboration patterns and underlying reasons. It will be valuable to expand the scope to commercial tools for more comprehensive understanding.
@@ 3.1 Tool Collection
After a careful consideration, we compared human-AI collaboration patterns of the latest tools after Jun. 2023 with those of earlier tools to study the research trend of human-AI collaborative tool in the era of generative AI. First, the public release dates of two representative large-scale generative AI techniques, ChatGPT and Stable Diffusion, are Nov. 30, 2022111https://openai.com/index/chatgpt/ and Aug. 22, 2022222https://stability.ai/news/celebrating-one-year-of-stable-diffusion, respectively. Considering the delay led by tool development and paper publishing, the research tools to apply them and other similar techniques might only be available in mid-2023. Then, we examined the tools collected by Li et al. [21]. These tools were published by Jun. 2023 and did not have extensive application of these models. Our collected tools further verify the decision: there were pioneering ones that apply text-to-image models [41] and LLMs [34, 30, 24] published in VIS 2023 in Oct. 2023.
@@
Table 1: This table presents the human-AI collaboration patterns and the applied generative AI techniques in our collected tools. The collaborators’ roles are shown in abbreviation: H-C, H-A, H-O, H-R are the creator, assistant, optimizer, and reviewer roles performed by humans; A-C, A-A, A-O, A-R are the corresponding roles by AI. The techniques are denoted with the icons: * refers to LLMs and # refers to text-to-image models.
Tool Analysis Planning Implementation Communication
[42] N/A N/A H-C N/A
[46] N/A N/A H-C A-A ∗# N/A
[6] N/A N/A H-C A-A ∗# N/A
[2] N/A N/A H-C A-A N/A
[31] N/A N/A H-C A-A N/A
[3] N/A N/A H-C A-A N/A
[34] N/A N/A A-C H-O N/A
[36] N/A H-C H-C A-A N/A
[38] N/A H-C H-C A-A N/A
[10] N/A H-C H-C H-C A-A
[30] N/A H-C A-C N/A
[27] N/A H-C A-C H-O N/A
[41] H-C A-A N/A H-C A-A # N/A
[24] H-C N/A A-C H-O N/A
[37] H-C H-C A-A A-C H-O N/A
[45] H-C A-A H-C A-C N/A
[19] H-C A-A H-C A-A A-C N/A
[28] A-C A-C A-C N/A
[32] A-C A-C A-C N/A
[43] A-C A-C A-C A-C
[18] A-C A-C A-C A-C
[1] A-C A-C A-C A-C
[5] A-C H-O A-C H-O A-C H-O N/A
[16] A-C H-O A-C H-O A-C H-O N/A
[7] A-C H-O A-C H-O A-C H-O N/A
[40] A-C H-R A-C H-R A-C N/A
[11] N/A N/A N/A H-C A-A
@@
Figure 1: This figure shows the change of human-AI collaboration pattern frequencies between the latest and earlier tools in four stages. In a subfigure, the color of blocks and numbers in the heatmap encodes the change of pattern frequencies in the stage. A pattern is represented by a human role on the horizontal axis and an AI role on the vertical axis. Green and positive numbers mean that the pattern is more frequent in the latest tools while red and negative numbers indicate that the pattern is less frequent. A number in brown means that the pattern is newly identified in the latest tools. An empty white cell means that the pattern has not emerged in any tools.
3.2 Pattern Coding
@@
After collecting all tools, we coded their human-AI collaboration patterns with the framework proposed by Li et al. [21]. The framework has two dimensions: the roles of human and AI collaborators (i.e., human-creator and AI-creator to complete the majority of tasks in each stage, human-assistant and AI-assistant to help creators on designated tasks in a stage, human-optimizer and AI-optimizer to improve the quality of the output of one task or the entire stage autonomously, and human-reviewer and AI-reviewer to provide suggestions for the output from a task or stage) and their collaboration stages (i.e., analysis stage to derive data insights and knowledge from datasets, planning stage to decide how to convey analytical findings through a data story, implementation stage to create data story content following the plan, and communication stage to share data stories directly with audiences). One author coded all tools after understanding the coding principles and the coded tools by Li et al. comprehensively. During the coding process, the results were also calibrated with previously coded tools as references. The final results are presented in Table 1 and an online browser: https://human-ai.notion.site/data-storytelling-dec-2024. In Table 1, to help readers compare the tools in two papers more conveniently, the order of tools aligns with the sorting method used in Li et al. [21], which is based on the stages each tool covers. In each stage, the leading role (i.e., creator) in one stage will be shown first. To facilitate comparison, we computed and visualized the changes in the relative frequencies [^3] All frequencies refer to relative frequencies to accommodate the different numbers of tools in each stage in earlier and the latest tools. of collaboration patterns in Fig. 1. The frequency of a pattern $p$ in a stage $s$ is computed as $f=Occurrence_{p\textrm{ in }s}/NumTool_{s}$ and the frequency change between the latest tools and earlier tools is $f_{latest}-f_{earlier}$. In Fig. 1, a positive change means that the pattern is gaining more attention while a negative one shows the frequency is declining. Our supplementary material also present the relative frequency of collaboration patterns among earlier and the latest tools for our readers to understand what patterns are widely studied.
We applied the framework for three-fold reasons. First, the framework distinguishes the collaboration patterns of human and AI at a fine granularity. Furthermore, the framework was summarized based on an interview study [20] that envisions potential AI application. Evaluating the alignment between existing tools and user expectations can help us reflect on current practices and identify future opportunities. Another advantage is that the framework is technique-agnostic, allowing us to compare collaboration patterns without interference led by differences in techniques.
@@
Combining the results of ours and Li et al. [21], we notice a considerable growth in the number of data storytelling tools in 2024. Compared to the 11 tools published in 2023, the 20 tools in 2024 nearly double the previous number, reaching a historic high. Furthermore, 16 out of the 27 latest tools apply generative AI models (Table 1). These models are most frequently used in the implementation stage (15/27), following up by the analysis (5/27), the planning (4/27), and the communication stages (3/27). The findings reveal that large-scale generative AI models has been widely applied across the storytelling workflow, supporting our motivation.
4 Findings
This section presents insights into the evolving collaboration patterns studied within the community, including consistently widely studied patterns, newly explored patterns, and those gaining increased attention. It is important to note that changes in patterns should not be interpreted as indicators of success or failure. They only reflect shifts in research interest.
4.1 Which Patterns Are Still Widely Studied?